If the circle x2 + y2 = a2 cuts off a chord of length 2b from the line y = mx + c , then (1) (1 -m2) (a2 + b2) = c2 (2) (1 -m2) (a2 -b2) = c2 (3) (1 + m2)(a2-b2) = c2

Grade:11

1 Answers

pradeep
5 years ago
If the circle x2 + y2 = a2 cuts off a chord of length 2b from the line y = mx + c , then (1) (1 -m2) (a2 + b2) = c2 (2) (1 -m2) (a2 -b2) = c2 (3) (1 + m2)(a2-b2) = c2